A Bibliography of Publications of Jorge Moré

Jorge Moré

Mathematics and Computer Science Division Argonne National Laboratory, 9700 S. Cass Ave.,

Argonne, IL 60439-4844 USA

Tel: +1 708 252 7238 FAX: +1 708 252 5986

E-mail: more@mcs.anl.gov (Internet)

14 June 2019 Version 1.15

Abstract

This bibliography records publications of Jorge Moré.

Title word cross-reference

n [MR73].

-dimensional [MR73].

1 [GHM80, MGH80].

2 [ACM91, AM91, ACMX92].

Active [BM88]. Algorithm [CGM84a, CGM85a, MC80, MGH81a, Mor78]. Algorithms [MT89, Mor83, Mor91, MT92]. application [DM74]. architectures [AM93]. automatic [AMB^{+}94].

bound [Mor88a, MT89, Mor91, MT91].

Bounds [CM87b]. BRENTM [MC80]. Broyden [MT76].

C5 [MGH81a]. characterization [DM74]. Classes [Mor74a, MR73]. Cody [MM09]. Coercivity [Mor74b]. collection [ACM91, AM91, ACMX92, Mor90]. coloring [CM83, CM84]. Complementarity [Mor74b, Mor74a]. Computing [AMB^{+}94, MS83]. concave [MV91]. Conditions [Mor74b, Mor74a]. constrained [CM87a, Mor88a, MT89, Mor91]. Constraints [BM88, BMT90, BM92, MT91]. Convergence [BMT90, BDM73, DM74, MT76]. convex [BMT90]. curvature [MS79].

decrease [MT92]. design [Mor80]. developments [Mor83]. differentiation [AMB^{+}94]. dimensional [MR73]. directions [MS79].

E4 [MGH81a]. Equations [MC79, MC80]. Estimating [CGM84a, CGM84b, CGM85b, CGM85a]. Estimation [CM83, CM84]. Evaluation [AM93]. Exposing [BM92].

F5 [MC80]. feasibility [Mor74a]. FORTRAN [CGM85a, MGH81a, CGM84a, MC80]. function [Mor74a]. functions [MR73].

Generalizations [Mor93]. global [MT76]. gradient [CM87a]. gradients [Mor88b]. graph [CM83, CM84]. guaranteed [MT92]. Guide [AM91, GHM80, MGH80, MW93].

Hessian [CM84, CGM85a, CGM85b].

Identification [BM88]. Implementation [GHM80, Mor79, Mor78].

Jacobian [AMB^{+}94, CM83, CGM84a, CGM84b]. Jim [MM09].

knapsack [MV91].

large [AM93, AMB^{+}94, Mor91, MT91]. large-scale [AM93, Mor91]. Levenberg [Mor78]. libraries [MM09]. Line [MT92]. linear [BMT90]. linearly [CM87a]. local [BDM73].

mappings [MR73]. Marquardt [Mor78]. mathematical [MM09]. Matrices [CGM84a, CGM84b, CGM85b, AMB^{+}94, CM83, CM84, CGM85a]. Method [MS84, MT76, MS79]. Methods [DM77, BDM73, BMT90, CM87a, DM74, Mor83]. MINPACK [ACM91, AM91, ACMX92, GHM80, MGH80, MSHG84]. MINPACK-1 [GHM80, MGH80]. MINPACK-2 [ACM91, AM91, ACMX92]. model [Mor90]. modified [MS79]. Motivation [DM77]. negative [MS79]. Newton [MS84, BDM73, CM87b, DM74, DM77, MS79]. Nonlinear [Mor74b, MC79, MC80, Mor74a, Mor90]. Notes [Mor82]. Numerical [MC79, MC80, Mor88a].

Optimization [MGH81a, MGH81b, MW93, AM93, Mor79, Mor80, Mor82].

P [MR73]. P- [MR73]. parallel [AM93]. performance [Mor91]. Pioneer [MM09]. Preliminary [ACM91]. problem [ACM91, AM91, ACMX92, Mor93]. Problems [Mor74b, AM93, CM87a, CM83, CM84, Mor74a, Mor88a, MT89, Mor90, Mor91, MV91, MT91]. programming [MT89, MT91]. Project [MSHG84]. Projected [CM87a, Mor88b]. properties [BMT90].

quadratic [MT89, MT91]. Quasi [CM87b, DM77, BDM73, DM74]. Quasi-Newton [CM87b, DM77, BDM73, DM74].

region [BMT90, MS83, Mor83, Mor93]. regions [Mor88b]. related [MR73].

S [MR73]. S-functions [MR73]. scale [AM93, Mor91]. search [MT92]. Software [CGM84b, CGM85b, MGH81a, MGH81b, MW93, Mor79, Mor80, Mor82, Mor83, MM09]. Solution [MC79, MC80, Mor88a, MV91, MT91]. Sparse [CGM84a, CGM84b, CGM85b, AMB^{+}94, CM83, CM84, CGM85a]. step [MS83]. Subroutine [MC80]. Subroutines [CGM84a, MGH81a, CGM85a]. sufficient [MT92]. superlinear [BDM73, DM74].

test [ACM91, AM91, ACMX92]. Testing [MGH81a, MGH81b, Mor79]. Theory [DM77, Mor78]. Trust [Mor88b, BMT90, MS83, Mor83, Mor93].

Unconstrained [MGH81a, MGH81b]. Updates [CM87b]. use [MS79]. User [AM91, MGH80]. using [AMB^{+}94].

vector [AM93]. Version [ACM91].

References

Averick:1991:MTP

[ACM91] B. M. Averick, R. G. Carter, and J. J. Moré. The MINPACK-2 test problem collection (Preliminary Version). Technical Memorandum ANL/MCS-TM-150, Argonne National Laboratory, Argonne, IL, USA, 1991. Also issued as Preprint 91-76 of the Army High Performance Computing Research Center at the University of Minnesota.

Averick:1992:MTP

[ACMX92] B. M. Averick, R. G. Carter, J. J. Moré, and G-L. Xue. The MINPACK-2 test problem collection. Preprint MCS-P153-0692, Argonne National Laboratory, Argonne, IL, USA, 1992.

Averick:1991:UGM

[AM91] B. M. Averick and J. J. Moré.

User guide for the MINPACK-2 test problem collection. Technical Memorandum ANL/MCS-TM-157, Argonne National Laboratory, Argonne, IL, USA, 1991. Also issued as Preprint 91-101 of the Army High Performance Computing Research Center at the University of Minnesota.

Averick:1993:ELS

[AM93] B. M. Averick and J. J. Moré. Evaluation of large-scale optimization problems on vector and parallel architectures. Preprint MCS-P379-0893, Argonne National Laboratory, Argonne, IL, USA, 1993.

Averick:1994:CLS

[AMB^{+}94] B. M. Averick, J. J. Moré, C. H. Bischof, A. Carle, and Andreas Griewank. Computing large sparse Jacobian matrices using automatic differentiation. SIAM J. Sci. Stat. Comput., 15:285–294, 1994. CODEN SIJCD4. ISSN 0196-5204.

Broyden:1973:LSC

[BDM73] C. G. Broyden, J. E. Dennis, and J. J. Moré. On the local and superlinear convergence of quasi-Newton methods. J. Inst. Math. Appl., 12:223–245, 1973.

Burke:1988:IAC

[BM88] James V. Burke and Jorge J. Moré. On the identification of active constraints. SIAM Journal on Numerical Analysis, 25(5): 1197–1211, October 1988. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Burke:1992:EC

[BM92] J. V. Burke and J. J. Moré. Exposing constraints. Preprint MCS-P308-0592, Argonne National Laboratory, Argonne, IL, USA, 1992.

Burke:1990:CPT

[BMT90] J. V. Burke, J. J. Moré, and G. Toraldo. Convergence properties of trust region methods for linear and convex constraints. Math. Programming, 47:305–336, 1990. CODEN MHPGA4. ISSN 0025-5610.

Coleman:1984:AFS

[CGM84a] Thomas F. Coleman, Burton S.

Garbow, and Jorge J. Moré. Algorithm 618: Fortran subroutines for estimating sparse Jacobian matrices. ACM Trans. Math. Software, 10(3):346–347, September 1984. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Coleman:1984:SES

[CGM84b] Thomas F. Coleman, Burton S.

Garbow, and Jorge J. Moré. Software for estimating sparse Jacobian matrices. ACM Trans. Math. Software, 10(3):329–345, September 1984. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Coleman:1985:AFS

[CGM85a] Thomas F. Coleman, Burton S.

Garbow, and Jorge J. Moré. Algorithm 636: FORTRAN subroutines for estimating sparse Hessian matrices. ACM Trans. Math. Software, 11(4):378, December 1985. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (elec-

tronic). URL http://www.acm.
org/pubs/citations/journals/
toms/1985-11-4/p378-coleman/

. The title of this paper incorrectly said Algorithm 649; it should be Algorithm 636.

Coleman:1985:SES

[CGM85b] Thomas F. Coleman, Burton S. Garbow, and Jorge J. Moré. Software for estimating sparse Hessian matrices. ACM Trans. Math. Software, 11(4):363–377, December 1985. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (elec-

                                        tronic). URL http://www.acm.
                                        org/pubs/citations/journals/
                                        toms/1985-11-4/p363-coleman/

.

Coleman:1983:ESJ

[CM83] Thomas F. Coleman and Jorge J. Moré. Estimation of sparse Jacobian matrices and graph coloring problems. SIAM Journal on Numerical Analysis, 20(1):187– 209, February 1983. CODEN SJ-NAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Coleman:1984:ESH

[CM84] T. F. Coleman and J. J. Moré. Estimation of sparse Hessian matrices and graph coloring problems. Math. Programming, 28:243–270, 1984. CODEN MHPGA4. ISSN 0025-5610.

Calamai:1987:PGM

[CM87a] P. H. Calamai and J. J. Moré. Projected gradient methods for linearly constrained problems. Math. Programming, 39:93–116, 1987. CODEN MHPGA4. ISSN 0025-5610.

Calamai:1987:QNU

[CM87b] Paul H. Calamai and Jorge J.

Moré. Quasi-Newton updates with bounds. SIAM Journal on Numerical Analysis, 24(6):1434– 1441, December 1987. CO-DEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Dennis:1974:CSC

[DM74] J. E. Dennis and J. J. Moré.

A characterization of superlinear convergence and its application to quasi-Newton methods. Math. Comp., 28:549–560, 1974.

Dennis:1977:QNM

[DM77] J. E. Dennis, Jr. and Jorge J.

Moré. Quasi-Newton methods, motivation and theory. SIAM Review, 19(1):46–89, January 1977. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Garbow:1980:IGM

[GHM80] B. S. Garbow, K. E. Hillstrom, and J. J. Moré. Implementation guide for MINPACK-1. Report ANL-80-68, Argonne National Laboratory, Argonne, IL, USA, 1980.

More:1979:NSN

[MC79] Jorge J. Moré and Michel Y.

Cosnard. Numerical solution of nonlinear equations. ACM Trans. Math. Software, 5(1):64– 85, March 1979. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

More:1980:ABF

[MC80] J. J. Moré and M. Y. Cosnard. Algorithm 554: BRENTM, A Fortran subroutine for the numerical solution of nonlinear equations [F5]. ACM Trans. Math. Software, 6(2):240–251, June 1980. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

More:1980:UGM

[MGH80] J. J. Moré, B. S. Garbow, and K. E. Hillstrom. User guide for MINPACK-1. Report ANL-80-74, Argonne National Laboratory, Argonne, IL, USA, 1980.

More:1981:AFS

[MGH81a] J. J. Moré, B. S. Garbow, and K. E. Hillstrom. Algorithm 566: FORTRAN subroutines for testing unconstrained optimization software [C5 [E4]]. ACM Trans. Math. Software, 7(1):136– 140, March 1981. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See also [?].

More:1981:TUO

[MGH81b] Jorge J. Moré, Burton S. Garbow, and Kenneth E. Hillstrom. Testing unconstrained optimization software. ACM Trans. Math. Software, 7(1):17–41, March 1981. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

More:2009:JCP

[MM09] Jorge Moré and Cleve Moler. Jim Cody: Pioneer in mathematical software libraries. NA Digest postings, June 28, 2009. URL http:

//www.netlib.org/na-digest-
html/09/v09n26.html#1.

More:1974:CFF

[Mor74a] J. J. Moré. Classes of function and feasibility conditions in nonlinear complementarity problems. Math. Programming, 6:327– 338, 1974. CODEN MHPGA4. ISSN 0025-5610.

More:1974:CCN

[Mor74b] Jorge J. Moré. Coercivity conditions in nonlinear complementarity problems. SIAM Review, 16 (1):1–16, ???? 1974. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

More:1978:LMA

[Mor78] J. J. Moré. The Levenberg– Marquardt algorithm: Implementation and theory. In G. A. Watson, editor, Numerical Analysis: Proceedings of the Biennial Conference held at Dundee, June 28-July 1, 1977 (Lecture Notes in Mathematics #630), pages 104– 116 (of xii + 199). Springer-Verlag, Berlin, Heidelberg, New York, Tokyo, 1978. LCCN QA1 .L471 v.630.

More:1979:ITO

[Mor79] J. J. Moré. Implementation and testing of optimization software. In L. D. Fosdick, editor, Performance Evaluation of Numerical Software, pages 253–266. North-Holland, Amsterdam, The Netherlands, 1979. ISBN 0-444-85330-8. LCCN QA297 .I18 1978.

More:1980:DOS

[Mor80] J. J. Moré. On the design of optimization software. Quad. Unione Mat. Italiana, 17:85–102, 1980.

More:1982:NOS

[Mor82] J. J. Moré. Notes on optimization software. In M. J. D. Powell, editor, Nonlinear Optimization 1981, pages 339–352 (of xvii + 559). Academic Press, New York, NY, USA, 1982. ISBN 0-12-563860-4. LCCN QA402.5 .N66 1981.

More:1983:RDA

[Mor83] J. J. Moré. Recent developments in algorithms and software for trust region methods. In A. Bachem, M. Gr¨ otschel, and B. Korte, editors, Mathematical Programming Bonn 1982 - The State of the Art, pages 258–287 (of viii + 655). Springer-Verlag, Berlin, Heidelberg, New York, Tokyo, 1983. ISBN 0-387-12082-3. LCCN QA402.5 .M3529 1983.

More:1988:NSB

[Mor88a] J. J. Moré. Numerical solution of bound constrained problems. In J. Noye and C. Fletcher, editors, Computational Techniques and Applications: CTAC-87, pages 29–37 (of xix + 685). North-Holland, Amsterdam, The Netherlands, 1988. ISBN 0-444-70400-0. LCCN QA299.6 .I561 1987.

More:1988:TRP

[Mor88b] J. J. Moré. Trust regions and projected gradients. In M. Iri and K. Yajima, editors, Systems Modelling and Optimization: Proceedings of the 13th IFIP Conference, Tokyo, Japan, August 31-September 4, 1987, Lecture Notes in Control and Information Sciences 113, pages 1–13 (of x + 787). Springer-Verlag, Berlin, Heidelberg, New York, Tokyo, 1988. ISBN 0-387-19238-7. LCCN QA402 .S95823 1988.

More:1990:CNM

[Mor90] J. J. Moré. A collection of nonlinear model problems. In E. L. Allgower and K. Georg, editors, Computational Solution of Nonlinear Systems of Equations, volume 26 of Lecture Notes in Applied Mathematics, pages 723–762 (of xix + 762). Amer. Math. Soc., Providence, RI, USA, 1990. ISBN 0-8218-1131-2. LCCN QA372 .C6374 1990.

More:1991:PAL

[Mor91] J. J. Moré. On the performance of algorithms for large-scale bound constrained problems. In T. F. Coleman and Y. Li, editors, Large-Scale Numerical Optimization, pages 31–45 (of 255). SIAM, Philadelphia, PA, USA, 1991. ISBN 0-89871-268-8. LCCN QA402.5 .L35 1990.

More:1993:GTR

[Mor93] J. J. Moré. Generalizations of the trust region problem. Optimization Methods and Software, 2:189– 209, 1993. ISSN 1055-6788.

More:1973:PFR

[MR73] J. J. Moré and W. C. Rheinboldt.

On P- and S-functions and related classes of n-dimensional mappings. Linear Algebra and Its Applications, 6:45–68, 1973.

More:1979:UDN

[MS79] J. J. Moré and D. C. Sorensen. On the use of directions of negative curvature in a modified Newton method. Math. Programming, 16: 1–20, 1979. CODEN MHPGA4. ISSN 0025-5610.

More:1983:CTR

[MS83] J. J. Moré and D. C. Sorensen. Computing a trust region step. SIAM J. Sci. Stat. Comput., 4: 553–572, 1983. CODEN SIJCD4. ISSN 0196-5204.

More:1984:NM

[MS84] J. J. Moré and D. C. Sorensen. Newton’s method. In G. H. Golub, editor, Studies in Numerical Analysis, pages 29–82 (of x + 415). Mathematical Association of America, Washington, DC, 1984. ISBN 0-88385-126-1. LCCN QA297 .S83 1984.

More:1984:MP

[MSHG84] J. J. Moré, D. C. Sorensen, K. E. Hillstrom, and B. S. Garbow. The MINPACK project. In W. J. Cowell, editor, Sources and Development of Mathematical Software, pages 88–111 (of xii + 404). Prentice-Hall, Englewood Cliffs, N.J., 1984. ISBN 0-13-823501-5. LCCN QA76.95 .S68 1984.

More:1976:GCB

[MT76] J. J. Moré and J. A. Trangenstein. On the global convergence of Broyden’s method. Math. Comp., 30: 523–540, 1976.

More:1989:ABC

[MT89] J. J. Moré and G. Toraldo. Algorithms for bound constrained quadratic programming problems. Num. Math, 55:377–400, 1989. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

More:1991:SLQ

[MT91] J. J. Moré and G. Toraldo. On the solution of large quadratic programming problems with bound constraints. SIAM J. Optimization, 1:93–113, 1991. CODEN SJOPE8. ISSN 1052-6234 (print), 1095-7189 (electronic).

More:1992:LSA

[MT92] J. J. Moré and D. J. Thuente. Line search algorithms with guaranteed sufficient decrease. Preprint MCS-P330-1092, Argonne National Laboratory, Argonne, IL, USA, 1992.

More:1991:SCK

[MV91] J. J. Moré and S. Vavasis. On the solution of concave knapsack problems. Math. Programming, 49:397–441, 1991. CODEN MH-PGA4. ISSN 0025-5610.

More:1993:OSG

[MW93] J. J. Moré and Stephen J.

Wright. Optimization Software Guide. SIAM, Philadelphia, PA, USA, 1993. ISBN 0-89871-322-6. xii + 154 pp. LCCN QA402.5 .M67 1993.